OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 2x + 9y = –30; (–8, 0) y = 2/9x + 36 y = 9/2x + 36 y = -9/2x + 36 y = 2/9x – 8

OpenStudy (anonymous):

@sebaxtiangb

OpenStudy (anonymous):

@MathleticMagicain

OpenStudy (anonymous):

@ShadowLegendX:

OpenStudy (shadowlegendx):

Do you know how to start this?

OpenStudy (anonymous):

nope

OpenStudy (shadowlegendx):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 2x + 9y = –30; (–8, 0) First youʻre going to have to solve for the line 2x + 9y - 30 to get the correct slope. Then find the line that is perpendicular to that slope reverse, negative it Then, we will use point slope form to input in those points, and our perpendicular line, to find our answer

OpenStudy (shadowlegendx):

I have to go soon, so I will try to get this displayed asap

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

I also have a few more

OpenStudy (shadowlegendx):

First we simplify 2x + 9y = - 30 Basically, we want to get 2x + 9y = - 30 into a y = mx + b format with m standing for slope and b standing for the y intercept(the point in which the line crosses the y axis) 2x + 9y = - 30 Subtract 2x from both sides 9y = -2x - 30 Divide the entire equation by 9 y = -2/9x - 30/9

OpenStudy (shadowlegendx):

Now we have our slope, -2/9x We reverse it to -9/2x and then change the sign to positive 9/2x

OpenStudy (anonymous):

D?

OpenStudy (shadowlegendx):

From there we go to point slope form y - y1 = m(x - x1) (–8, 0) -8 is the x 0 is the y y - 0 = m(x --8) y - 0 = m(x + 8) Input in our new slope y - 0 = 9/2(x + 8) Now we simplify First get rid of that 0, and then distribute y = 9/2x + 36 That is our answer

OpenStudy (anonymous):

5.5 units right y = | x | – 5.5 y = | x – 5.5 | y = | x + 5.5 | y = | x | + 5.5

OpenStudy (shadowlegendx):

I gtg, Iʻll be back on in a couple hours, I have to endure the traffic on the way back home :P

OpenStudy (anonymous):

What time will you be on tomorrow?

OpenStudy (shadowlegendx):

mhm, I have school, so...like 10, 11hrs from now? When Iʻm free after school, I gtg now, bye

OpenStudy (anonymous):

ok. tomorrow night at this time?